if an aerosol can contains a gas at 1.52atm pressure at 22°c, what will be the gas pressure in the can if it is dumped into an incinerator at 935°c?
1
Expert's answer
2019-10-17T07:18:37-0400

Initial pressure of can"(P_1)=1.52 \ atm"

Initial temperature of can"(T_1)=22\degree C=22+273=295K"

Final temperature of Can"(T_2)=935\degree C=935+273=1208K"

The volume of can remains constant.

Using the equation,

"\frac{P_1}{T_1}=\frac{P_2}{T_2}"

Substituting the values,

"\frac{1.52}{295}=\frac{P_2}{1208}"

"P_2=6.22\ atm"